# Georgy Lukyanov Toulouse School of Economics Anna Vlasova

### v1

@@ 1 Introduction
The model yields testable predictions. Holding fundamentals fixed, highly reputed experts recommend the risky action less frequently but achieve higher success rates when they do, because implementers work harder. Transitory reputation shocks (e.g., early successes) reduce the frequency of risky calls but make subsequent failures more revealing and more reputationally costly; negative shocks have the opposite effect. These predictions echo empirical patterns in settings where effort is observable or proxied (e.g., adherence) and where performance differences across experts are documented (Haskard Zolnierek and DiMatteo, 2009; Birkhäuer et al., 2017; Birkmeyer et al., 2003; Birkmeyer et al., 2013). They speak to the design of advisory relationships in medicine (operate vs. conservative care), finance (buy vs. hold and trade size), innovation management (greenlight vs. delay), and public policy (reform vs. status quo), where the scale and diligence of implementation are endogenous to the expert’s standing. [^1] A practical recipe for constructing reputation and effort proxies is in Online Appendix LABEL:OA-oa:measurement. Computational details for value iteration, cutoff computation, and simulations are in Online Appendix LABEL:OA-oa:alg.
@@
The paper proceeds as follows. Section 2 reviews related work. Section 3 sets up the environment, belief updates, and recursive equilibrium. Section 4 presents the threshold equilibrium, a binary–signal illustration, reputational conservatism, comparative statics, and reputation dynamics with a learning–absorption decomposition. Section 5 treats the surgery application and policy design via success–contingent bonuses, and covers monitoring, committees, endogenous exit, and a continuous–time approximation. Core proofs are in Appendices AD; algorithms, Gaussian formulas, robustness, committee variants, behavioral twists, and policy calibration appear in the Online Appendix LABEL:OA-oa:measurementLABEL:OA-oa:policy).
@@ 2 Related Literature
A second strand studies strategic expert advice when advisers care about reputation. In reputational cheap talk with continuous signals, experts pool messages to manage reputational risk, yielding coarse disclosure (Ottaviani and Sørensen, 2006). We provide a dynamic counterpart with a state-dependent cutoff that moves with public reputation. Conformity and caution under career concerns have been emphasized in settings where “being wrong alone” is worse than “being wrong with the crowd” (Scharfstein and Stein, 1990; Prendergast, 1993; Morris, 2001). In our model, similar caution emerges through the effort channel: when clients work harder for highly reputed experts, a failure is more revealing, pushing such experts toward conservatism. With short-lived customers, initial honesty may profitably build a reputation that can later be exploited (Ozyurt, 2016). Paradoxically, very competent advisers with spotless records may be fired because their behavior is inferred to be uninformative or biased (Schottmüller, 2019); our law of motion clarifies when occasional visible failures are actually informative (and thus desirable) versus reputation-damaging. Relatedly, “good lies” can build credibility in the long run (Pavesi and Scotti, 2017); in our environment, explicit outcomes discipline such strategies because failures become more diagnostic when reputation is high.
@@ 3.4 Agent’s best response
with $e^{*}(a,\pi)$ increasing in $\lambda(a,\pi)$. Under the common quadratic benchmark $c(x)=\tfrac{1}{2}x^{2}$, one has $e^{*}(a,\pi)=\lambda(a,\pi)$. [^5] General primitives for $c(\cdot)$ (strictly convex $C^{2}$ and effort caps) and robustness to outcome/observability frictions (baseline success under $a=0$, misclassification, partial observability of $a$) are developed in Online Appendix LABEL:OA-oa:robustness.
@@ 4 Core Results
This section establishes the equilibrium structure and the main comparative statics. Throughout we maintain Assumptions A1–A4 from Section 3. Proofs that are longer or purely technical are deferred to the appendices as indicated. [^7] Algorithms and replication details for all figures and numerical objects appear in Online Appendix LABEL:OA-oa:alg.
@@ 4.2 Cutoff structure in the private signal
Proof (short). Monotonicity: For any bounded $V$, the map $\pi\mapsto\phi\,\rho(\pi)+\delta\,\mathbb{E}[V(\pi^{\prime})\mid\pi]$ is increasing because posteriors are increasing in prior odds under MLRP (A1) and $V$ is evaluated at larger posteriors in the first-order stochastic dominance sense. Contraction holds with modulus $\delta$ in the sup norm. For convexity, fix a cutoff policy $s(\cdot)$. Under $a=1$, the posterior takes $\{\pi^{+}(\pi),\pi^{-}(\pi)\}$ with probabilities that shift toward the failure branch as reputation (and thus effort) rises; Condition (6) ensures the induced posterior kernel becomes a mean-preserving spread in the convex order as $\pi$ increases. Hence $\pi\mapsto\mathbb{E}[V(\pi^{\prime})\mid\pi]$ is convex for convex $V$, and so is $T^{\,s}V$. The fixed point $V^{\,s}$ is therefore increasing and convex; the equilibrium value $V=V^{\,s^{*}}$ inherits the properties. Full details appear in Appendix LABEL:OA-app:A-convex. ∎
@@ 5 Discussion and Extensions
This section develops several extensions that either sharpen the main mechanisms or broaden applicability. [^8] Persistent environments and correlated information are treated in Online Appendix LABEL:OA-oa:persistence. In each case we preserve the unified notation from Sections 34 and keep the public state one-dimensional (the reputation $\pi$). Formal proofs and additional details are provided in the appendices referenced below.
@@ 5.4 Multiple implementers and committees
The result quantifies a natural complementarity: reputational incentives are stronger when a risky recommendation triggers broader implementation capacity, because the chance of success—and hence the expected reputational gain—is higher. Appendix G provides the details and variants with success thresholds (e.g., $k$-out-of-$m$ success) and correlated efforts. [^11] General monotone aggregation rules, exchangeable/heterogeneous committees, and comparative statics beyond $k$-of-$n$ are in Online Appendix LABEL:OA-oa:committees.
@@ 5.5 Policy design: success-contingent bonuses
By strict single crossing, $s\mapsto\Delta_{H}(s;\pi)$ is strictly increasing; adding $\alpha\beta_{1}$ shifts $\Delta_{H}$ up and yields a unique $s^{*}_{\beta_{1}}(\pi)$, strictly decreasing and continuous in $\beta_{1}$. Since $\rho(\pi;\cdot)$ is a continuous, strictly decreasing function of the cutoff (with positive densities at $s^{*}$), it is continuous and strictly increasing in $\beta_{1}$. As $\beta_{1}\downarrow 0$, $s^{*}_{\beta_{1}}(\pi)\to s^{*}(\pi)$ and $\rho(\pi;\beta_{1})\to\rho(\pi;0)$; as $\beta_{1}\uparrow\infty$, $s^{*}_{\beta_{1}}(\pi)\to$-$\infty$ and $\rho(\pi;\beta_{1})\to 1$. The affine case follows by replacing $\alpha\beta_{1}$ with $\alpha\beta_{1}-(1-\alpha)\beta_{0}$. Full details are in Online Appendix LABEL:OA-oa:policy.
@@ 5.8 Policy and design: contingent fees in dynamic settings
Thus, modest explicit incentives can partially undo reputational conservatism and restore experimentation at high reputation. [^12] Planner’s welfare and the first-order condition linking $\partial\rho/\partial\beta$ to per-experiment surplus and the budget shadow price, as well as affine $(\beta_{1},-\beta_{0})$ contracts (including corners and implementation sets), are presented in Online Appendix LABEL:OA-oa:policy. See also the minimal-implementation formulas in (31).

### v2

@@ 1 Introduction
The model yields testable predictions. Holding fundamentals fixed, highly reputed experts recommend the risky action less frequently but achieve higher success rates when they do, because implementers work harder. Transitory reputation shocks (e.g., early successes) reduce the frequency of risky calls but make subsequent failures more revealing and more reputationally costly; negative shocks have the opposite effect. These predictions echo empirical patterns in settings where effort is observable or proxied (e.g., adherence) and where performance differences across experts are documented (Haskard Zolnierek and DiMatteo, 2009; Birkhäuer et al., 2017; Birkmeyer et al., 2003; Birkmeyer et al., 2013). They speak to the design of advisory relationships in medicine (operate vs. conservative care), finance (buy vs. hold and trade size), innovation management (greenlight vs. delay), and public policy (reform vs. status quo), where the scale and diligence of implementation are endogenous to the expert’s standing. [^1] A practical recipe for constructing reputation and effort proxies is in Online Appendix OA.2. Computational details for value iteration, cutoff computation, and simulations are in Online Appendix OA.3.
@@
The paper proceeds as follows. Section 2 reviews related work. Section 3 sets up the environment, belief updates, and recursive equilibrium. Section 4 presents the threshold equilibrium, a binary–signal illustration, reputational conservatism, comparative statics, and reputation dynamics with a learning–absorption decomposition. Section 5 treats the surgery application and policy design via success–contingent bonuses, and covers monitoring, committees, endogenous exit, and a continuous–time approximation. Core proofs are in Appendices AD; algorithms, Gaussian formulas, robustness, committee variants, behavioral twists, and policy calibration appear in the Online Appendix OA.2OA.8).
@@ 2 Related Literature
A second strand studies strategic expert advice when advisers care about reputation. In reputational cheap talk with continuous signals, experts pool messages to manage reputational risk, yielding coarse disclosure (Ottaviani and Sørensen, 2006). We provide a dynamic counterpart with a state-dependent cutoff that moves with public reputation. Conformity and caution under career concerns have been emphasized in settings where “being wrong alone” is worse than “being wrong with the crowd” (Scharfstein and Stein, 1990; Prendergast, 1993; Morris, 2001). In our model, similar caution emerges through the effort channel: when clients work harder for highly reputed experts, a failure is more revealing, pushing such experts toward conservatism. With short-lived customers, initial honesty may profitably build a reputation that can later be exploited (Ozyurt, 2016). Paradoxically, very competent advisers with spotless records may be fired because their behavior is inferred to be uninformative or biased (Schottmüller, 2019); our law of motion clarifies when occasional visible failures are actually informative (and thus desirable) versus reputation-damaging. Relatedly, “good lies” can build credibility in the long run (Pavesi and Scotti, 2022); in our environment, explicit outcomes discipline such strategies because failures become more diagnostic when reputation is high.
@@ 3.4 Agent’s best response
with $e^{*}(a,\pi)$ increasing in $\lambda(a,\pi)$. Under the common quadratic benchmark $c(x)=\tfrac{1}{2}x^{2}$, one has $e^{*}(a,\pi)=\lambda(a,\pi)$. [^5] General primitives for $c(\cdot)$ (strictly convex $C^{2}$ and effort caps) and robustness to outcome/observability frictions (baseline success under $a=0$, misclassification, partial observability of $a$) are developed in Online Appendix OA.4.
@@ 4 Core Results
This section establishes the equilibrium structure and the main comparative statics. Throughout we maintain Assumptions A1–A4 from Section 3. Proofs that are longer or purely technical are deferred to the appendices as indicated. [^7] Algorithms and replication details for all figures and numerical objects appear in Online Appendix OA.3.
@@ 4.2 Cutoff structure in the private signal
Proof (short). Monotonicity: For any bounded $V$, the map $\pi\mapsto\phi\,\rho(\pi)+\delta\,\mathbb{E}[V(\pi^{\prime})\mid\pi]$ is increasing because posteriors are increasing in prior odds under MLRP (A1) and $V$ is evaluated at larger posteriors in the first-order stochastic dominance sense. Contraction holds with modulus $\delta$ in the sup norm. For convexity, fix a cutoff policy $s(\cdot)$. Under $a=1$, the posterior takes $\{\pi^{+}(\pi),\pi^{-}(\pi)\}$ with probabilities that shift toward the failure branch as reputation (and thus effort) rises; Condition (6) ensures the induced posterior kernel becomes a mean-preserving spread in the convex order as $\pi$ increases. Hence $\pi\mapsto\mathbb{E}[V(\pi^{\prime})\mid\pi]$ is convex for convex $V$, and so is $T^{\,s}V$. The fixed point $V^{\,s}$ is therefore increasing and convex; the equilibrium value $V=V^{\,s^{*}}$ inherits the properties. Full details appear in Appendix OA.1. ∎
@@ 5 Discussion and Extensions
This section develops several extensions that either sharpen the main mechanisms or broaden applicability. [^8] Persistent environments and correlated information are treated in Online Appendix OA.5. In each case we preserve the unified notation from Sections 34 and keep the public state one-dimensional (the reputation $\pi$). Formal proofs and additional details are provided in the appendices referenced below.
@@ 5.4 Multiple implementers and committees
The result quantifies a natural complementarity: reputational incentives are stronger when a risky recommendation triggers broader implementation capacity, because the chance of success—and hence the expected reputational gain—is higher. Appendix G provides the details and variants with success thresholds (e.g., $k$-out-of-$m$ success) and correlated efforts. [^11] General monotone aggregation rules, exchangeable/heterogeneous committees, and comparative statics beyond $k$-of-$n$ are in Online Appendix OA.6.
@@ 5.5 Policy design: success-contingent bonuses
By strict single crossing, $s\mapsto\Delta_{H}(s;\pi)$ is strictly increasing; adding $\alpha\beta_{1}$ shifts $\Delta_{H}$ up and yields a unique $s^{*}_{\beta_{1}}(\pi)$, strictly decreasing and continuous in $\beta_{1}$. Since $\rho(\pi;\cdot)$ is a continuous, strictly decreasing function of the cutoff (with positive densities at $s^{*}$), it is continuous and strictly increasing in $\beta_{1}$. As $\beta_{1}\downarrow 0$, $s^{*}_{\beta_{1}}(\pi)\to s^{*}(\pi)$ and $\rho(\pi;\beta_{1})\to\rho(\pi;0)$; as $\beta_{1}\uparrow\infty$, $s^{*}_{\beta_{1}}(\pi)\to$-$\infty$ and $\rho(\pi;\beta_{1})\to 1$. The affine case follows by replacing $\alpha\beta_{1}$ with $\alpha\beta_{1}-(1-\alpha)\beta_{0}$. Full details are in Online Appendix OA.8.
@@ 5.8 Policy and design: contingent fees in dynamic settings
Thus, modest explicit incentives can partially undo reputational conservatism and restore experimentation at high reputation. [^12] Planner’s welfare and the first-order condition linking $\partial\rho/\partial\beta$ to per-experiment surplus and the budget shadow price, as well as affine $(\beta_{1},-\beta_{0})$ contracts (including corners and implementation sets), are presented in Online Appendix OA.8. See also the minimal-implementation formulas in (31).